# HEDN: A Hard-Easy Dual Network with Source Reliability Assessment for Cross-Subject EEG Emotion Recognition

### v2

@@ IV-D2 Evaluation Results on DEAP dataset
Table III summarizes the performance of HEDN on the DEAP dataset. We compare HEDN with several representative baselines, including DANN, MSDA-SFE, and PR-PL, which are widely used in EEG emotion recognition. ATDD-LSTM and DISD-Net are also included, as they are specifically optimized for the DEAP dataset. In addition, DBPM is incorporated as the current state-of-the-art method. Overall, HEDN achieves superior performance across both the valence and arousal dimensions, reaching 73.20% and 73.22%, respectively. Notably, HEDN maintains consistently strong performance even when compared with models that have been specially optimized for DEAP (ATDD-LSTM, DISD-Net), demonstrating its generalizability and confirming its robustness and effectiveness in addressing cross-subject variability.
@@ IV-E Cross-Dataset Evaluation
Methods $\text{SEED}\to\text{SEED-IV}$ $\text{SEED-IV}\to\text{SEED}$
@@ V-A Ablation Study
For the architectural and objective-level components of HEDN, we examine six aspects. (1) w/o Hard Network removes the Hard Network entirely, leading to a substantial accuracy drop of more than 10% on both datasets, since the Hard Network is essential for mitigating the distribution discrepancy between source and target domains. (2) w/o Easy Network removes the Easy Network, resulting in the accuracy on the SEED dataset decreasing to 85.52%, which highlights its critical role in label propagation. (3) w/o $\mathcal{L}_{clu}$ on the Source, (4) w/o $\mathcal{L}_{clu}$ on the Target, and (5) w/o $\mathcal{L}_{clu}$ on the Source and Target correspond to the ablation of Eq. (13). The results show that removing $\mathcal{L}_{clu}$ entirely yields the largest accuracy degradation, and $\mathcal{L}_{clu}$ on the target domain exhibits a greater influence on the SEED dataset. Moreover, $\mathcal{L}_{clu}$ is evidently more crucial on SEED than on SEED-IV; on SEED-IV, removing it only reduces the accuracy from 79.36% to 77.85%, yielding a small degradation of 1.51%, which suggests that the cluster structure in SEED-IV is more stable and less sensitive to training perturbations. (6) w/o Step-Wise Joint Optimization, which optimizes all objectives in a single step, causes even greater performance degradation than removing the second stage (i.e., w/o $\mathcal{L}_{clu}$ on both the Source and Target), indicating that the two-step optimization scheme, which directs the model to focus on distinct tasks at different stages, helps maintain stable cluster structures and facilitates reliable label propagation.
@@ V-B Complexity Analysis
To evaluate the efficiency of the proposed HEDN model, we compare it with several publicly available baselines from Table II that provide open-source implementations. As shown in Table VI, HEDN achieves high accuracy while substantially improving adaptation efficiency. Specifically, HEDN requires only 37.54 seconds for adaptation, which is more than four times faster than the next-best competitor and up to 200 times faster than the slowest baseline. This efficiency derives from its SRA dynamic routing: at each iteration, HEDN selects only the most and least reliable source domains and routes them to two specialized branches. By selectively processing informative samples rather than uniformly traversing all source data, HEDN eliminates redundant computations, maximizes information gain per update, and reduces overall computational burden.
Although the per-iteration SRA evaluation increases the FLOPs count (93.72M) compared to other methods, HEDN converges rapidly and remains compact, with only 36,932 parameters and a model size of 219.08 KB. Its inference latency of 7.93 ms is well within acceptable limits for real-time EEG processing, particularly considering the substantial improvements in both adaptation speed and accuracy. Consequently, HEDN achieves an optimal balance between efficiency and performance: it avoids the high adaptation costs of complex multi-source methods (e.g., MSMDA, S2A2-MSDA) while outperforming single-source approaches (e.g., DANN, PR-PL) in both speed and accuracy.
@@ V-E2 Convergence Stability
As shown in Table VIII, HEDN achieves the highest final accuracy on the SEED dataset (84.17%) with a moderate gap of $\Delta_{\text{acc}}=10.02\%$, demonstrating that it not only attains state-of-the-art peak performance but also sustains a high level of accuracy at convergence. On the more challenging SEED-IV dataset, HEDN exhibits a larger gap $\Delta_{\text{acc}}=20.32\%$, but its final accuracy (59.04%) remains competitive, second only to S2A2-MSDA (61.43%) and MSMDA (59.71%).
@@ V-F Source Number Evaluation
From an efficiency perspective, both computational cost and adaptation time increase with the number of sources. The FLOPs grow from 24.05M with 2 sources to 93.72M with 14 sources, and the total adaptation time increases from 20.34 seconds to 37.54 seconds. Nevertheless, even in the full 14-source setting, HEDN remains computationally efficient, requiring fewer than 100M FLOPs and less than 40 seconds for adaptation on the SEED dataset. These results demonstrate that HEDN offers a practical and scalable solution for real-world EEG emotion recognition. Users can adjust the number of source domains to balance accuracy and resource constraints, making the framework adaptable to diverse deployment scenarios—from lightweight edge devices to high-performance systems.
@@ V-G Impact of Density-Based Prototype Initialization
Initialization Method SEED SEED-IV
@@
DBSCAN (Silhouette Tuning) 92.19 76.20

### v3

@@ IV-D2 Evaluation Results on DEAP dataset
Table III summarizes the performance of HEDN on the DEAP dataset. We compare HEDN with several representative baselines, including DANN, MSDA-SFE, and PR-PL, which are widely used in EEG emotion recognition. ATDD-LSTM and DISD-Net are also included, as they are specifically optimized for the DEAP dataset. In addition, DBPM is incorporated as the current state-of-the-art method. Overall, HEDN achieves superior performance across both the valence and arousal dimensions, reaching 73.07% and 73.20%, respectively. Notably, HEDN maintains consistently strong performance even when compared with models that have been specially optimized for DEAP (ATDD-LSTM, DISD-Net), demonstrating its generalizability and confirming its robustness and effectiveness in addressing cross-subject variability.
@@ IV-E Cross-Dataset Evaluation
Methods $\text{SEED}\to\text{SEED-IV}$ (%) $\text{SEED-IV}\to\text{SEED}$ (%)
@@ V-A Ablation Study
For the architectural and objective-level components of HEDN, we examine six aspects. (1) w/o Hard Network removes the Hard Network entirely, leading to a substantial accuracy drop of more than 10% on both datasets, since the Hard Network is essential for mitigating the distribution discrepancy between source and target domains. (2) w/o Easy Network removes the Easy Network, resulting in the accuracy on the SEED dataset decreasing to 85.19%, which highlights its critical role in label propagation. (3) w/o $\mathcal{L}_{clu}$ on the Source, (4) w/o $\mathcal{L}_{clu}$ on the Target, and (5) w/o $\mathcal{L}_{clu}$ on the Source and Target correspond to the ablation of Eq. (13). The results show that removing $\mathcal{L}_{clu}$ entirely yields the largest accuracy degradation, and $\mathcal{L}_{clu}$ on the target domain exhibits a greater influence on the SEED dataset. Moreover, $\mathcal{L}_{clu}$ is evidently more crucial on SEED than on SEED-IV; on SEED-IV, removing it only reduces the accuracy from 79.36% to 78.81%, yielding a small degradation of 0.55%, which suggests that the cluster structure in SEED-IV is more stable and less sensitive to training perturbations. (6) w/o Step-Wise Joint Optimization, which optimizes all objectives in a single step, causes even greater performance degradation than removing the second stage (i.e., w/o $\mathcal{L}_{clu}$ on both the Source and Target) on the SEED-IV dataset, indicating that the two-step optimization scheme, which directs the model to focus on distinct tasks at different stages, helps maintain stable cluster structures and facilitates reliable label propagation.
@@ V-B Complexity Analysis
To evaluate the efficiency of the proposed HEDN model, we compare it with several publicly available baselines from Table II that provide open-source implementations. As shown in Table VI, HEDN achieves high accuracy while substantially improving adaptation efficiency. Specifically, HEDN requires only 48.98 seconds for adaptation, which is more than four times faster than the next-best competitor and up to 160 times faster than the slowest baseline. This efficiency derives from its SRA dynamic routing: at each iteration, HEDN selects only the most and least reliable source domains and routes them to two specialized branches. By selectively processing informative samples rather than uniformly traversing all source data, HEDN eliminates redundant computations, maximizes information gain per update, and reduces overall computational burden.
Although the per-iteration SRA evaluation increases the FLOPs count (61.55M) compared to other methods, HEDN converges rapidly and remains compact, with only 36,932 parameters and a model size of 219.14 KB. Its inference latency of 9.72 ms is well within acceptable limits for real-time EEG processing, particularly considering the substantial improvements in both adaptation speed and accuracy. Consequently, HEDN achieves an optimal balance between efficiency and performance: it avoids the high adaptation costs of complex multi-source methods (e.g., MSMDA, S2A2-MSDA) while outperforming single-source approaches (e.g., DANN, PR-PL) in both speed and accuracy.
@@ V-E2 Convergence Stability
As shown in Table VIII, HEDN achieves the highest final accuracy on the SEED dataset (84.17%) with a moderate gap of $\Delta_{\text{acc}}=9.83\%$, demonstrating that it not only attains state-of-the-art peak performance but also sustains a high level of accuracy at convergence. On the more challenging SEED-IV dataset, HEDN exhibits a larger gap $\Delta_{\text{acc}}=20.32\%$, but its final accuracy (59.04%) remains competitive, second only to S2A2-MSDA (61.43%) and MSMDA (59.71%).
@@ V-F Source Number Evaluation
From an efficiency perspective, both computational cost and adaptation time increase with the number of sources. The FLOPs grow from 19.45M with 2 sources to 61.55M with 14 sources, and the total adaptation time increases from 22.76 seconds to 48.98 seconds. Nevertheless, even in the full 14-source setting, HEDN remains computationally efficient, requiring fewer than 65M FLOPs and less than 50 seconds for adaptation on the SEED dataset. These results demonstrate that HEDN offers a practical and scalable solution for real-world EEG emotion recognition. Users can adjust the number of source domains to balance accuracy and resource constraints, making the framework adaptable to diverse deployment scenarios—from lightweight edge devices to high-performance systems.
@@ V-G Impact of Density-Based Prototype Initialization
Initialization Method SEED (%) SEED-IV (%)
@@
DBSCAN (Silhouette Tuning) 92.19 78.48